### v2

# Barren-plateau free variational quantum simulation of $\mathbb{Z}_{2}$ lattice gauge theories

Abstract
In this work, we design a variational quantum eigensolver (VQE) suitable for investigating ground states and static string breaking in a $\mathbb{Z}_{2}$ lattice gauge theory (LGT). We consider a two-leg ladder lattice coupled to Kogut-Susskind staggered fermions and verify the results of the VQE simulations using tensor network methods. We find that for varying Hamiltonian parameter regimes and in the presence of external charges, the VQE is able to arrive at the gauge-invariant ground state without explicitly enforcing gauge invariance through penalty terms. Additionally, experiments showing string breaking are performed on IBM’s quantum platform. Thus, VQEs are seen to be a promising tool for $\mathbb{Z}_{2}$ LGTs, and could pave the way for studies of other gauge groups. We find that the scaling of gradients with the number of qubits is favorable for avoiding barren plateaus. At the same time, it is not clear how to efficiently simulate the LGT using classical methods. Furthermore, strategies that avoid barren plateaus arise naturally as features of LGTs, such as choosing the initialization by setting the Gauss law sector and restricting the Hilbert space to the gauge-invariant subspace.
I Introduction
Gauge theories, i.e. field theories with a local symmetry, are the theoretical foundation for the description of particle physics. At the same time, analytical access in the non-perturbative regime is often not possible. One way forward is lattice gauge theory (LGT), in which space-time is discretized on a lattice [1]. In this framework, matter fields are assigned to the lattice sites, while gauge fields are associated with the connecting links, all while maintaining gauge symmetry. In particular, this discretization allows for a numerical approach with sophisticated Monte Carlo methods.
While Monte Carlo methods have demonstrated excellent predictive power for static problems, they are limited by the sign problem in certain parameter regimes, leaving many relevant questions unanswered. This has triggered the ongoing effort to develop quantum and quantum-inspired methods for LGTs, such as tensor networks (TNs), to overcome these limitations. Current quantum processors are a long way from being used for realistic simulations of gauge theories. So far, successful demonstrations have been carried out for low-dimensional toy models [2, 3, 4, 5, 6, 7] of confinement [8, 9, 10], and scattering dynamics  [11, 12, 13, 14, 15]. For a review, see  [16, 17, 18].
In this work, we study a $\mathbb{Z}_{2}$ LGT [19] on a two-leg ladder geometry coupled to Kogut-Susskind fermions [20]. The two-leg ladder is an intermediate step between one and two dimensions. It is the simplest geometry that features magnetic plaquette interactions, and therefore all of the terms of the LGT Hamiltonian, while still being efficiently verifiable using tensor network techniques. The simplicity of the $\mathbb{Z}_{2}$ LGT is amenable to noisy intermediate-scale quantum (NISQ) devices, while still exhibiting complex phenomena, such as confinement.
Confinement is the mechanism that explains why isolated quarks or gluons are not observed in nature. The dynamical process of string breaking is fundamental to confining gauge theories, and is described by quantum chromodynamics. Using $\mathbb{Z}_{2}$ LGT as a toy model, we investigate static aspects of string breaking as a consequence of confinement. When two charges are confined in the $\mathbb{Z}_{2}$ LGT, by Gauss law they are necessarily connected by an electric flux tube. The flux tube experiences a linearly rising potential proportional to their distance. From a certain distance it becomes energetically favorable to break the flux string, and here this is done by forming an additional pair of charges.
We apply a variational quantum eigensolver (VQE) to $\mathbb{Z}_{2}$ LGT on the two-leg ladder lattice to recover ground state properties and to explore static string breaking. VQEs are hybrid quantum-classical algorithms for estimating expectation values. Such variational algorithms are widely studied in the NISQ era as they do not require error-corrected qubits. However, they have no provable guarantees, so the question remains: are there problems that variational quantum algorithms are better suited to solve compared to classical methods.
While we do not yet have the quantum hardware to answer this question, recent investigations of variational quantum algorithms have unearthed many issues. In particular, there seems to be a connection between the trainability of a VQE (the absence of barren plateaus) and efficient classical simulation [21]. Thus, it is important that proposals involving barren-plateau free VQEs consider the scaling with system size and the effects of noise.
In this work, we find features within the framework of $\mathbb{Z}_{2}$ LGT that are compatible with common strategies for avoiding barren plateaus, all the while being nontrivial to classically simulate. Warm starts are one such strategy, where one chooses to initialize close to a solution, where there may be narrow gorges [22]. Analogously, for investigations of LGT, it is necessary to initialize to some Gauss law sector, which thereby restricts the search landscape. The search process for a VQE is benefited by constraining to an explorable subspace of Hilbert space [23]; by gauge symmetry for LGTs, this is the subspace of the gauge-invariant, physical, states.
We draw upon these features using two types of ansatz circuits. One is restricted entirely to the symmetric subspace of physical states by being composed of gauge invariant terms. The other has overlap with states outside of the physical subspace, but nonetheless arrives at gauge invariant solutions over the course of the optimization. For the simulations we consider scaling with system size and numbers of shots, to find good convergence of ground states and we are able to witness static string breaking. Results are also obtained using IBM’s quantum hardware, where we perform experiments in the region of the transition into the string-broken phase. These findings underscore the usefulness of VQEs in the investigation of LGTs.
II Background and methods
II.1 Lattice gauge theories
Figure 1: Features of LGTs on the two-leg ladder lattice. The LGT specifies matter sites and links between them pertaining to the underlying gauge field. The arrows on the links show the orientation of the lattice. We have labeled $\hat{E}_{x}$ and $\hat{E}_{y}$ the orientations of the electric field and the ring labeled $\hat{P}$ indicates the magnetic plaquette term. For the vertex $\bm{l}$, we have indicated the notation for the adjacent link and matter site that is one unit vector $\hat{x}$ across. The red arrows show how we linearly arrange the qubits to perform the Jordan-Wigner transformation: we count sites from the leftmost lower corner, then traverse upwards, then rightwards along the lattice. The alternating colors of the matter sites illustrate the coupling to staggered fermions where the sign alternates. On this diagram we have also included an example string configuration, through the solid arrow connecting two charges that have been placed. The overlapping circles show the change in sign to indicating the placed charges.
The defining feature of gauge theories is invariance under local transformations. This leads to the presence of a local conservation law or a continuity equation that determines the interplay between the matter and gauge degrees of freedom. In the Hamiltonian formulation using temporal gauge, this results in an additional constraint known as Gauss law.
Lattice gauge theories (LGTs) provide a means to probe gauge theories numerically. Numerical studies such as those using Monte Carlo techniques are limited by the sign problem in certain parameter regimes, thus a long-term goal for quantum computing is to access these intractable regions. One way to discretize gauge theory considers placing matter sites on the vertices and the underlying gauge group acts on the connecting links.
The pure gauge part of the theory is constructed following Ref. [19] and the lattice notation for the two-dimensional ladder geometry we consider is illustrated in Fig. 1. When the underlying gauge group is $\mathbb{Z}_{2}$, the unitary operators acting on the links fulfill the $\mathbb{Z}_{2}$ algebra, so they can be re-written in terms of the Pauli matrices $X$ and $Z$. For the ladder geometry, the resulting pure gauge part of the Hamiltonian is given by:
[…]
where the operators act on the links emanating from the matter site with coordinate $\bm{l}=(l_{x},l_{y})\in\mathbb{Z}^{2}$, and $\hat{x}$, $\hat{y}$ are unit vectors in $x$ and $y$-direction. The first part of this equation is the electric energy, where $X$ takes the role of an electric field. The second term corresponds to the magnetic energy contribution and represents a product of the ZZ operators on a plaquette of the lattice. The parameter $\mu>0$ tunes the relative strength between the electric and magnetic energy.
To couple the theory to fermionic matter, we consider Kogut-Susskind staggered fermions [20]. This formulation essentially corresponds to separating the spinor components to different lattice sites, resulting in alternating the signs of the masses and charges across the lattice as reflected in the orange and blue coloring in Fig. 1.
Then the terms representing the gauge-matter interaction and the fermion mass read:
$H_{m}$ $=J\sideset{}{{}_{\bm{l}}}{\sum}\left(\phi_{\bm{l}}^{\dagger}Z_{\bm{l},\bm{l}+\hat{k}}\phi_{\bm{l}+\hat{k}}+\text{h.c.}\right)$ (2)
$+m\sideset{}{{}_{\bm{l}}}{\sum}(-1)^{l_{1}+l_{2}}\phi^{\dagger}_{\bm{l}}\phi_{\bm{l}},$
where $J,m>0$, $\phi_{\bm{l}}$ is the spinless fermionic operator at site $\bm{l}$, and the connections $Z_{\bm{l},\bm{l}+\hat{k}}$ lift the gauge symmetry to be local for this Hamiltonian, that otherwise consists of nearest-neighbor interactions. In the limit $J\rightarrow 0$ the gauge and matter fields do not interact, hence we refer to it as the non-interacting limit.
[…]
$G_{\bm{l}}=(-1)^{l_{1}+l_{2}}(-1)^{n_{l}}\sideset{}{{}_{\hat{k}}}{\prod}X_{\bm{l},\bm{l}+\hat{k}}\sideset{}{{}_{\hat{k}}}{\prod}X_{\bm{l}-\hat{k}^{\prime},\bm{l}}.$ (3)
The prefactor represents the different nature of the odd and even lattice sites, and we define the number operator $n_{l}=\phi_{l}^{\dagger}\phi_{l}$. Depending on the point of the lattice, the operators acting on the links are replaced by the boundary condition $X_{\bm{l},\bm{l}+\hat{x}}\!\rightarrow 1$.
Since $\forall\bm{l}$, $H,G_{\bm{l}}]=0$, $G_{\bm{l}}$ is a local symmetry that decomposes the Hilbert space into separate sectors depending on the charge. Taking into account the staggering, we can consider the Gauss law sector of the states we perform calculations with.
For simulations of LGTs, it is necessary for physical results to be gauge invariant, making it necessary to enforce Gauss law. One way to enforce gauge invariance is through penalty terms that are proportional to (3). Gauge-invariant states subject to gauge-invariant operators remain in this physical subspace, i.e., the symmetry remains intact. This subspace is still exponentially large, while being non-trivial to split into further gauge-invariant subspaces  [24, 25].
Incorporating terms proportional to (3) is also a means to introduce infinitely massive static charges to the theory. A consequence of Gauss law is that a pair of charges placed on matter sites in vicinity of each other are necessarily connected by a flux tube or string. These charges experience a linearly rising potential that is proportional to their separation.
From a certain distance onward, it becomes energetically favorable to break the string, which creates an additional pair of charges. This is known as static string breaking, and is depicted in Fig. 1, where a configuration of a string is shown between two sites with charges, indicated by the opposite coloring. Static string breaking and confinement has also been investigated for $\mathbb{Z}_{2}$ LGT in the following works [26, 27, 28, 10].
Then, the total Hamiltonian is as follows,
$H=H_{g}+H_{m}+V\sideset{}{{}_{\bm{l}}}{\sum}(G_{\bm{l}}-q_{\bm{l}})^{\dagger}(G_{\bm{l}}-q_{\bm{l}}),$ (4)
with $V>0$. Placing charges amounts to adding penalty terms on the appropriate site with the opposite sign.
To compute expectation values of this Hamiltonian on a quantum computer, we perform the Jordan-Wigner transformation [29] on the matter Hamiltonian (2) that couples to fermionic degrees of freedom. The gauge part and penalty terms already consist of Pauli operators. For this we must choose a linear ordering of the fermions. In Fig. 1, the ordering zig-zags from the bottom leg to the top, running across the rungs towards the left, as indicated by the red arrows. Subsequently, we translate the fermions to spins using $\phi_{\bm{l}}=\sideset{}{{}_{k<l}}{\prod}(iZ_{k})\sigma_{l}^{-}$, with $\sigma^{\pm}=\frac{1}{2}(X\pm iY)$ and $l$ being the linear index of the ordering of the fermionic fields as described.
II.2 Variational quantum eigensolver
[…]
For observables formed of tensor products of simple Pauli operators, the quantum processor can efficiently evaluate the expectation value. A classical optimizer then improves on this guess of parameters $\bm{\theta}$, thus completing this quantum-classical loop.
The choice of ansatz must carefully balance incorporating physical information of the problem, i.e., being expressible, with being trainable. Expressibility is the extent to which the ensemble of unitaries generated by the ansatz contains the unitary that is ‘close’ to that which minimizes the cost. Then for classical gradient-based optimizers, trainability means that the cost landscape should manifest large enough gradients to optimize the parameters and minimize the cost function. Barren plateaus occur when the gradients of the cost function (5) become exponentially suppressed with the system size. Thus the features we require of the VQE seem to be at odds – deeper circuits are in general more expressible, but by construction are susceptible to barren plateaus.
Circumstantially, there seems to be the additional issue that the trainability of a VQE implies that it is classically simulable. We can avoid barren plateaus by using ansatz circuits that explore smaller regions of Hilbert space related to the underlying problem structure. One way to reduce the dimensionality of the parameter space is to apply the symmetries of the system. Herein lies what enables efficient classical simulation – this reduction tends to identify some polynomially large subspace [21].
As such, it is possible to diagnose barren plateaus from the size of the dynamical Lie algebra (DLA) spanned by the gate generators. For the gates of the parametrized ansatz circuit, the DLA generates any of the arbitrary circuits that it can express. In [30] it is conjectured that the gradients of the cost function scale inverse-polynomially with the dimension of the DLA. However, while a small DLA may avoid barren plateaus, a polynomial-sized DLA implies the problem is classically simulable [31]. Classical simulation can be performed using techniques such as low-weight Pauli propagation [32, 33].
These are important considerations for ansatz design. A choice of ansatz that considers the available native gates, we call a hardware efficient ansatz (HEA). Whereas an ansatz based on unitaries formed of the problem Hamiltonian is known as the Hamiltonian variational ansatz (HVA). For the local Hamiltonian $H=H_{K}$, where each $H_{K}$ consists of commuting Pauli operators, the HVA can be written [34, 35] as:
[…]
Initialization strategies to avoid barren plateaus have been proposed for both the HEA [36] and HVA [37]. However, such warm starts and initialization strategies show promise for VQEs, but are not guaranteed to find the global minimum, nor is there a general initialization strategy in subspaces containing the solution. This occurs for only a vanishing fraction of problem instances [22].
In [38] it is proven that local Hamiltonians training finite or logarithmic local-depth circuits (FLDC) will not encounter barren plateaus. Here the two-dimensional Toric code is considered, which is a variant of the pure gauge part of the $\mathbb{Z}_{2}$ LGT. In general FLDC cannot be classically simulated for two dimensions and higher.
VQEs are considered promising NISQ era algorithms since they do not require error correction, however they have no provable guarantees. To ask what problems can test VQEs at scale, invokes questions of trainability, so we know more and more the limitations of VQEs. It is then crucial to identify problems that are average-case hard to solve classically while being barren-plateau free.
Figure 2: The ansatz circuits considered in this work. These are circuits for the simplest example of one plaquette. The red gates set the initial state, while the green gates are parametrized, and form the repeating layers. Here we use the shorthand $X\rightarrow R_{X}(\theta)$, so each green gate has associated one variable angle. a) shows the gauge-invariant (GI) ansatz and b) the hardware-efficient, MBQC-inspired ($Z$) ansatz.
II.3 Tensor network states
[…]
In the expression above, the $M_{k}^{i_{k}}$ are complex $\chi\times\chi$ matrices for $1<k<N$ and $M_{1}^{i_{1}}$ ($M_{N}^{i_{N}}$) is a $\chi$-dimensional row (column) vector. Thus, for a given set of indices $i_{k}$, the coefficients of the wave function are parametrized by a product of matrices, hence the name of the ansatz. The matrix size $\chi$, commonly referred to as the bond dimension of the MPS, determines the number of parameters in the ansatz and limits the maximum amount of entanglement that can be present in the state.
MPS can be used as a family of variational ansätze to compute the ground state of a (local) Hamiltonian. To this end, the energy expectation value is variationally minimized by successively updating a tensor while keeping the others fixed. At each step, the optimal tensor is found by determining the ground state of an effective Hamiltonian describing the interaction of the site with its environment. Iterating this procedure until convergence, the final MPS provides an approximation for the ground state of the Hamiltonian. For our numerical calculations, we use the ITensor library where we run a standard two-site algorithm updating two neighboring tensors together [41, 42]. This allows for adaptively adjusting the bond dimension during the computation to obtain results with a certain precision. A detailed review on MPS methods can be found in Ref. [43].
Note that in order to apply the MPS approach to our ladder geometry, we have to map the local degrees of freedom sitting on the vertices and links to a linear chain. This does not create interactions whose range scales with the system size, so the resulting one-dimensional mapping can generally be addressed efficiently with MPS.
Figure 3: Scaling of the variance with system size across ansätze and numbers of layers. We consider the GI and ZZ ansatz, across 1, 2, and 3 layers, averaged over 100 samples. As the system size is increased, we see a favorable scaling in the variance, so that gradients remain non-vanishing when the number of qubits is beyond classical simulation.
III Results
In this section we present the findings of numerical experiments and experiments using IBM’s quantum hardware performing VQE. First we motivate the choice of ansatz circuits. We then consider questions of scalability and classical simulability. We go on to show results for ground state estimation and static string breaking, while considering fixed numbers of shots and the presence of noise. The numerical experiments are verified using tensor network calculations, and in doing this we find cases where the VQE can identify the ground state while the TN converges to local minima, so that those TN simulations require the addition of penalty terms.
III.1 Choice of ansatz
We consider two forms of ansatz circuits in this work. Fig. 2 shows the state initialization (red gates) and one of the parametrized repeating layers (green gates). For the simulations we are limited to three plaquettes (that amounts to 18 qubits) and use no more than three layers. In Fig. 2, qubits are assigned starting from the bottom left corner of the ladder (Fig. 1) and then traversing up then left. In this way, the qubit assignments run as matter- link- matter-qubits then link- link-qubits.
[…]
Enforcing gauge invariance can be difficult on a quantum computer since introducing large penalty terms can lead to barren plateaus. Preserving gauge invariance is also difficult, since errors in the system evolution can push the state outside of the physical Hilbert space.
One way to enforce and preserve gauge invariance is by constructing a gauge-preserving ansatz [44, 10], as shown in Fig. 2a. This ansatz, which we label GI, is a Hamiltonian variational ansatz, where the parametrized unitaries are each of the terms from the Hamiltonian Eq. (4). Upon transforming Eq. (2) to spin degrees of freedom using the Jordan-Wigner transformation acting on the relevant matter and link qubits, the kinetic term introduces the unitaries proportional to, e.g., $\exp{(-i\theta/2X\!\otimes\!Z\!\otimes\!Z\!\otimes\!Y)}$.
We also consider a simpler ansatz, that is presented in Fig. 2b and we refer to it as ZZ. The gates used in this ansatz are inspired by measurement-based quantum computing (MBQC) [45], consisting of single qubit rotation gates, and multi-qubit $Z$ rotations for the entangling layers. Such multi-qubit rotations can be implemented in MBQC by the introduction of one ancillary qubit [46]. The choice of simpler gates also makes this ansatz more hardware-efficient. The ansatz was designed by considering that the gauge part of the Hamiltonian in Eq. (1) has 4-body plaquette interactions and the kinetic part of the Hamiltonian in Eq. (2) introduces 3-body interactions along the edges connecting matter sites (see Fig. 1). We note that these are not truly 3-body terms as the transport law from gauge invariance involves intermediate matter qubits where the edge is not formed of adjacent qubits. These 3- and 4-body terms are accounted for through $R_{ZZZ}$ and $R_{ZZZZ}$ rotation gates on the relevant qubits. Furthermore this ansatz is gauge invariant when we initialize close to the angle $\theta\approx\pi$ for all the parametrized gates. Thus the ansatz has overlap with the gauge-invariant subspace and we can initialize the VQE in this space.
Implementing these hardware-efficient or Hamiltonian-variational ansätze alone would lead to slow convergence of the LGT VQE and trapping in local minima. For the GI ansatz, it is necessary to initialize in the correct Gauss law sector. This is done by setting the initial state appropriately, as seen in Fig. 2a, where we initialize to the zero-charge sector. The alternating $X$ gates take into account the staggered fermions. When charges are added to the theory, the Gauss law sector we initialize to, must be coordinated. Practically, this amounts to adding an extra $X$ gate to the matter qubit with the charge.
For the ZZ ansatz we can be more relaxed about the Gauss law sector, as can be seen in the simpler state initialization step in Fig. 2b. We ensure that we are beginning in some Gauss law sector by applying the Hadamard gate $H$ on the link qubits, putting those in a superposition. By initializing to these sectors, we see that LGTs provide a meaningful initialization strategy for the VQE.
Figure 4: Average fidelity of Gauss law operator $G_{l}$ across each site during the optimization. For the ZZ ansatz we plot how the fidelity varies for the Hamiltonian with $J=m=\mu=1$ over the course of the optimization using SLSQP. In dark green, we initialize in the gauge-invariant subspace by setting all the initial angles to $\pi$, whereas for light green this choice is random. When the optimization begins in this subspace, in a few iterations, the VQE explores states outside of the space to then return, arriving on the state that minimizes the cost function. When the optimization begins outside of this subspace, the VQE nonetheless converges upon the state within the physical subspace.
We find good properties of the choice of ansätze. In Fig. 3 we show how the variance of the gradient of the cost function scales with system size and numbers of layers for the choice of ansatz circuit. For the simulations we have averaged over 100 samples and find favorable scaling of the variance up to system sizes of four plaquettes and here the Hamiltonian is the same as in Fig. 5. We also find that for both ansatz circuits, when scaling from one to two plaquettes, computing the DLA has too large a computational cost. The GI ansatz has similarities with the DLA of the matchgate circuit, which is known to be barren-plateau free and classically simulable [47]. However, the presence of $X$ gates make it so that the size of the DLA for the GI ansatz of one plaquette is $4080$, while the matchgate circuit on the same number of qubits ($8$) has dimension $120$. Another such example where the scaling of the variance of the cost function changes upon the inclusion of additional unitaries is shown in [30].
Figure 5: Performance of the VQE. We variationally estimate the ground state of the Hamiltonian in Eq. (4) with $J=3$, $m=1$ and $\mu=2$, while increasing the number of plaquettes $P$, i.e., qubits NN, numbers of ansatz layers $L$, and number of shots. On the $y$-axis we plot the energy relative error ratio with the ground state energy $E_{GS}$ obtained using tensor network methods. As the system size is increased, the behavior of the VQE remains stable, which is a good indication for using the VQE for larger system sizes. For each instance we have taken the average of 2020 runs and used the SPSA optimizer.
In the noiseless case, low-weight Pauli propagation (PP) is a useful technique for computing expectation values where the problem is efficiently classically simulable [33]. However, for these simulations we find large errors and ineffective weight truncation. Importantly, these ansatz circuits use non-Clifford rotation gates, so that their entangling layers do not meet the criteria to use this method, but all the while evading barren-plateaus.
The LGT framework provides natural parallels with the features of VQEs that escape barren plateaus. For the average case it is not known how best to initialize in the cost landscape, while in LGT this is answered for by Gauss law sectors and the gauge invariant subspace. Gauss law also acts as a signal to when the simulation may be trapped in local minima, since Gauss law is violated for these states. For VQEs to work, the optimization uses symmetries of the problem, while here we have the gauge-invariant subspace. Furthermore, a problem is classically simulable if this subspace is polynomially sized, but we know the overall gauge-invariant subspace to be exponential in size [24, 25].
For the results, classical simulation of the VQE was performed using Pennylane [48], and to interact with IBM’s quantum platform, we have used Qiskit [49].
III.2 Ground state estimation
We explore using VQE to find the ground state over a wide range of parameter values, with each of $J$, $m$, $\mu$ equaling up to $20$. The limit $J\rightarrow 0$ is known as the non-interacting limit, and with stronger interactions $J$, we find that the ZZ ansatz performs less well. This performance is improved by including penalty terms, with $V\sim J$. We consider system sizes up to three plaquettes, and simulations with up to three repeating layers. The obtained ground state energy is compared to the tensor network (TN) calculation.
To better understand the way the VQE searches, we also consider the fidelity of the resultant state at each stage of the optimization with the Gauss law operator $G_{l}$ (3) to check for Gauss law violations, this is shown for an instance of the VQE in Fig. 4. This provides an idea of how the optimization explores the physical Hilbert space and its boundary. As we expect, the GI ansatz remains gauge invariant throughout the noiseless optimization, while for the ZZ ansatz, we would expect to require the enforcement of Gauss law through penalty terms. However, regardless of whether we initialize within the gauge-invariant space, the VQE will arrive at the gauge-invariant ground state.
We show results for varying the numbers of shots, system size and numbers of layers for an instance of a Hamiltonian configuration in Fig. 5. Here we use $J=3$, $m=1$, $\mu=2$, run simulations up to $3$ layers (see Fig. 2 for a schematic of an ansatz layer) and for system sizes up to $3$ plaquettes. We find that with increasing system size the performance of the VQE remains roughly constant. For lower numbers of shots, the ansatz does not perform better with increasing layers. We also see that one layer of the GI ansatz can obtain good convergence, so the ansatz is already quite expressive for this case.
In Fig. 6, we plot the optimization to find the ground state energy of the Hamiltonian in Eq. (4) with $J=5$, $m=1$, $\mu=2$ and charges placed at $(0,0)$ and $(0,2)$. This case is interesting as the corresponding TN calculation is stuck in a local minima, with this value plotted as a horizontal line. Thus the TN calculation for comparison requires the addition of penalty terms to the Hamiltonian, while this is not true for the VQE using either GI or ZZ ansatz. This can occur since TN approaches tend to be biased towards weakly correlated phases.
Figure 6: VQE optimization of the ground state. Here we consider the Hamiltonian in Eq. (4) with $J=5$, $m=1$, $\mu=2$ and charges placed at $(0,0)$ and $(2,0)$. The main curves show two instances of the full state-vector simulation for a single run of VQE optimized using SLSQP. The bold curve is a generic instance, while the translucent curve shows the case where the VQE escapes the local minima. The grey dashed line is the ground state energy obtained using TN methods and we can see the simulation is stuck in the same local minima. The black dashed line is the reference TN simulation with the addition of a penalty term $V=100$ that we compare our results with. Without the addition of penalty terms, the VQE is successful in converging to the ground state. In the inset we show results also for fixed numbers of shots $[500,1000,5000,10000]$ averaged over 1515 runs. In order to consider fixed numbers of shots, we optimize using SPSA. It seems the final energy does not depend on the number of shots. The black dashed line is the ground state obtained using tensor networks, while the blue dashed line is the average of the SLSQP runs. We note that on average, with this configuration SPSA attains values higher than if we optimized using SLSQP.
In the main plot of Fig. 6 we plot two instances of state vector simulation of the VQE with the GI ansatz. In one of these instances the simulation is able to converge towards the ground state energy, while the other is an example where the VQE escapes a local minimum – the local minima that the TN simulation converges to. In the subfigure, we plot results for varying numbers of shots $[500,1000,5000,10000]$, and we find that the behavior of the VQE is less affected by increasing the number of shots. We note here that experiments with fixed numbers of shots use the SPSA optimizer, which we find to be less performant than SLSQP in this setting.
III.3 String breaking
Figure 7: Observation of string breaking in the $\mathbb{Z}_{2}$ LGT on the $P=3$ plaquette ladder lattice. The topmost figure plots the static potential $V(d)$ against increasing lattice spacing $d$ between the charges. The main plot shows results for one run of VQE using the GI ansatz. The lattice spacing is measured from $(0,0)$, and one of the charges is kept placed here. Thus charges at $(0,1)$ and $(1,0)$ are $d=1$ apart, and for these equidistant pairs discernibility in their values of $V(d)$ is due to boundary effects. The average of the $V(d)$ is the black line, which shows the expected behavior of $V(d)$. We have a linearly increasing potential, up to the point where string breaking occurs and it becomes energetically favorable to break the string, thus $V(d)$ becomes constant. We also plot simulated values from one run using the ZZ ansatz. Here the optimization could have run for more iterations, however due to the compute time it was terminated here. The values of the ZZ ansatz are to be used in reference with the IBM results – this is described in the main text. Here we have also plotted the static potential of the state prepared on ibm_marrakesh using the angles from this $Z$ VQE simulation. The best result of five consecutive runs was chosen. In the bottom figure we depict configurations of the lattice at criticality, between $d=2$ and $d=3$. The expectation values $\langle P\rangle$ correspond to the observable $\langle Z\rangle$ on the matter sites and $\langle X\rangle$ along the links. With charges placed between $(0,0)$ and $(2,0)$ ($d=2$), we see the presence of a flux string, then extending this second charge to $(2,1)$ breaks the string, and we can see the values of $\langle Z\rangle$ becomes flipped, indicating charge generation.
We also use the VQE to investigate static string breaking. Here we consider the three plaquette lattice, where we place pairs of charges within the theory and increase the distance between them. In the Hamiltonian of Eq. (4), this surmounts to introducing site-dependent terms proportional to Eq. (3) that take the opposite sign in the coefficient. The notion of charge is thus charged or uncharged, rather than positive, negative, or uncharged. This can be observed for certain parameter regimes of Eq. (4), where the string breaking critical distance is on the order of the lattice size.
In Fig. 7 we present results for one run of the VQE for the Hamiltonian Eq. (4) with $J=5$, $m=1$, and $\mu=3$. The case $\mu=2$ is shown in the previous section. Between these values $\mu=2$ and $\mu=3$, there is a transition such that string breaking occurs at the lattice spacing $d=3$. For these simulations we perform full state vector simulation to focus on the string breaking transition. However, due to the previous results with fixed shots, we expect to still observe this transition here.
The results for string breaking are captured by the static potential $V(d)$, where $V(d)=E_{0}-E_{c}$, and $E_{c}$ is the ground state with charges in Fig. 7. The lattice spacing $d$ is taken with reference to the fermion at $(0,0)$, so $d$ counts the number of links between this and the other available fermion. This means that for $d=1,2,3$ there are two fermion sites and $d=4$ contains the furthest fermion from $(0,0)$ at $(3,1)$. The oscillations in $V(d)$ for the two fermions per fixed $d$ is due to the boundary conditions and superpositions of string configurations, where this effect is heightened at the critical $d=2$. Here the charge at $(2,0)$ still has the presence of a flux string, while for $(1,1)$ we find the string vanishes. The effects of left and right boundary conditions can be mitigated by considering longer ladder geometries, then studying the action of charges in the center. While for the superposition effects, where the flux string might take one of two paths, this choice can be imposed in the model.
The black line plots the average value of $V(d)$, where we see a initial linear increase in $V(d)$ as the charges are separated and the string grows longer, then for some $2<d<3$, it becomes energetically favorable to break the string by forming two new charges. This behavior is shown in the lattice configuration plots of Fig. 7. Here we have measured the observable $\langle Z\rangle$ on the matter qubits and $\langle X\rangle$ on the link qubits. The observable $\langle Z\rangle$ tells us the sign of the staggering and where $\langle X\rangle=1$ indicates flux. In the upper diagram, we see the presence of a string when a charge is placed at $(2,0)$. In the lower diagram, where a charge is placed at $(2,1)$, we see the effects of string breaking, as the flux string is no longer present, and the opposite sign of $\langle Z\rangle$ indicates charge creation.
Fig. 7 also contains results obtained using IBM’s 156 qubit Heron r2 processor. We demonstrate the state preparation required for a single iteration of the VQE using 3 layers of the ZZ ansatz. Running the entire VQE would be too costly, so we consider the state preparation step that is required at every iteration. The ZZ ansatz was chosen as it is more hardware-efficient. Using Qiskit we perform the VQE up to some fixed number of iterations. The state we arrive at is not fully converged yet as this would require prolonged compute time, but is gauge-invariant, so we use this. The resultant 141 angles are then used to prepare the state on ibm_marrakesh, where the transpiled circuit has a depth of 525 and a $CZ$ count of 436. We use the default IBM error mitigation of zero-noise extrapolation (ZNE) and Twirled Readout Error eXtinction (TREX) [50], and take a plug-and-play approach to the experiments, where we choose the best of 5 runs. For the complicated circuits it is remarkable that the qualitative behavior of the VQE can be seen, but the results indicate that we should consider smaller system sizes. The initial angles for state preparation can be improved by using small penalty terms with the ZZ ansatz, since this would lead to faster convergence so reduced compute time. This would also introduce another aspect of the experiments that requires fine-tuning, that is not possible with the limited access to hardware.
Figure 8: Observation of string breaking on quantum hardware. Here we consider the $P=2$ plaquette $\mathbb{Z}_{2}$ LGT with Hamiltonian parameters $J=0.5$, $m=1.5$, and $\mu=1$, which has a string breaking transition at $d=2$. We show results for the ground state energy with no charges ($d=0$) and pairs of charges with one fixed at $(0,0)$ and another at different lattice sites, with spacing $d$ from $(0,0)$ (see Fig. 7 for example lattice configurations). We plot the results for the TN simulation, one run of the state vector simulation of the VQE, and one run of state preparation using the optimized rotation angles from the VQE on ibm_marrakesh. The VQE is performed using three layers of the ZZ ansatz. The TN simulations and the VQE simulations have strong agreement, which means that the resulting angle parameters for the state preparation using IBM are very close to the ground state. The IBM results recover the qualitative behavior of string breaking and the two phases (existence of string and string-broken) can be distinguished by a nearly constant offset in the value of the energy arising from the errors.
Considering now the two plaquette LGT Hamiltonian, using TN simulations we find that with the parameters $J=0.5$, $m=1.5$, and $\mu=1$, there is a string breaking transition at $d=2$. We use this problem setting to explore string breaking using ibm_marrakesh. Again we consider state preparation using the quantum computer, so here we perform state vector simulation of the VQE using three layers of the ZZ ansatz, and then prepare the resulting state on the IBM device. The experiments are of a single run only (as though we are performing the last step of the VQE optimisation on the quantum device) and the results for the resulting ground state energy $E$ are plotted in Fig. 8, where we have used the default ZNE and TREX error mitigation.
For this case there is strong agreement of the VQE with three layers of the ZZ ansatz and the TN simulations. The smaller number of qubits makes it possible to reach convergence using less compute. This means that the resulting state we prepare on the quantum device is already close to the ground state. Considering only one run of the IBM state preparation, we immediately see that the qualitative behavior of string breaking is preserved. The oscillations at the critical $d=2$ are amplified on the noisy device. Charges at $d=2$ correspond to charges placed on the lattice at coordinates $(0,2)$ and $(1,1)$. Here we see the two regimes more clearly, where the charge placed at $(1,1)$ still has a flux string, whereas at $(0,2)$ the string is broken, so the energy is equal to the case with the charge at $(1,2)$ or $d=3$.
Again for these experiments we have taken a plug-and-play approach, so it is remarkable that these qualitative features can be seen. The two phases, being existence of string and string-broken, can also be distinguished by the size of the error. The energy that we converge to when there is a string at $d=1$ is much closer to the simulated value, while in the broken phase the offset is larger, yet constant. Thus the difference in how the errors manifest also qualitatively characterizes the two phases.
It seems that in the two plaquette case, the numbers of qubits and therefore numbers of interacting gates, is manageable enough for using IBM quantum hardware. The ZZ ansatz is also more hardware efficient, so it is interesting to test the limits of the hardware as errors accumulate. The GI ansatz has comparable performance with fewer numbers of layers, (see Fig. 5) so it would be interesting to perform similar analysis with one layer of the GI ansatz. The degree to which the ZZ ansatz is hardware efficient can also be determined in this way.
IV Discussion
In this work we use a VQE to investigate a $\mathbb{Z}_{2}$ LGT coupled to Kogut-Susskind fermions on a two-leg ladder lattice. We are able to recover ground state properties and aspects of static string breaking using this technique. In doing so, we find that $\mathbb{Z}_{2}$ LGTs have features that suit investigation using VQEs, since they avoid barren plateaus and can be nontrivial to classically simulate.
One way to avoid barren plateaus is to use smart initialization strategies [22]; however a general strategy to identify these or to avoid local minima is not known. For LGTs on the other hand, we always initialize in some Gauss law sector and would know beforehand which to initialize in. Further, checking for Gauss law violations can alert as to whether the solution is a local minimum.
Another way is to constrain the searchable Hilbert space. This is done through the choice of ansatz, which is commonly designed to respect the symmetries of the problem. Solutions of LGTs are constrained to the gauge-invariant subspace of Hilbert space, and here we present an ansatz of gauge-invariant Hamiltonian terms that remains in this space. When the subspace spanned by the ansatz is polynomial in size, the problem can be classically simulated. All the while, the full space of gauge-invariant states is exponentially large [24, 25].
In this work we present two ansatz circuits, one that is constructed of gauge-invariant Hamiltonian terms, this we refer to as GI, and another that is inspired by MBQC that we name ZZ, as it is composed of multi-qubit $Z$ rotation gates. The ZZ ansatz is remarkably effective for convergence of the VQE to the ground state, further, by tracking the fidelity with the Gauss law operator, we see how this ansatz explores Hilbert space, to arrive in the gauge-invariant subspace. Both of these ansatz demonstrate good scaling in the variance of the cost function with increasing system size. At the same time, there is evidence the DLA does not have some low polynomial power scaling with system size, indicating that it is not efficiently classically simulated.
Accordingly, LGTs seem to provide interesting problems for investigations using VQE. This $\mathbb{Z}_{2}$ LGT is trainable, while so far being non-trivial to efficiently classically simulate. We have also come across instances where the tensor network simulation is stuck in local minima, while the VQE approach is able to converge close to the ground state energy. This suggests that problems of LGTs could be useful for testing VQEs at scale.
Fundamentally, the VQE could never be performed in ideal conditions, so considering the effects of scaling with system size, number of shots and incorporating the effects of noise is vital for making such a determination. We find that even with low numbers of shots $\sim 500$, convergence to the ground state is close and the resulting state is gauge-invariant. We also find that when scaling with system size, this behavior remains constant, which indicates that for larger systems placed on quantum devices, the VQE could behave as we expect.
Preliminary studies of the effects of noise are performed using IBM’s quantum platform. Here we obtain results using the ZZ ansatz for a string breaking transition of a three plaquette $\mathbb{Z}_{2}$ LGT. Since these circuits decompose to become quite deep, we are surprised to find that with using the default error mitigation, and not fine-tuning the device, we arrive at low energy states. In a limited capacity, we explore this further, by performing experiments with a smaller two plaquette $\mathbb{Z}_{2}$ LGT. Here we are able to qualitatively distinguish the string breaking transition from only one run for each configuration on the quantum hardware. These results highlight the effectiveness of the VQE approach to $\mathbb{Z}_{2}$ LGT.
In future work we would look to perform more experiments using quantum hardware. For smaller system sizes, it would be useful to perform some rounds of optimization in the VQE to see whether the resulting state remains gauge-invariant. This can be extended to observations of the string-breaking transition, where it would be interesting to see if this persists. We would also be interested to see the effect of small penalty terms on the cost landscape, since small terms in a region $\sim J$ can benefit the convergence of the $ZZ$ ansatz. Here one can also investigate the effect of small penalty terms in mitigating noise that may push the state out of the gauge-invariant subspace. We would also consider larger ladder lattices to mitigate the boundary effects. Some preliminary studies involving square lattices have also been performed, and show that we can extend this approach.
The behavior of the GI ansatz using current hardware versus the more hardware-efficient ZZ ansatz can also be analyzed, which would indicate the practicality of scaling such a VQE in the NISQ era. For those studies it would be beneficial to work more closely with the quantum hardware provider. It would also be beneficial to extend these studies to other platforms too. For instance the long-range multi-qubit interaction gates of the two ansatz may better suit ion-trap [51] or neutral atom based [52] quantum hardware. It is also interesting to explore applying the ZZ ansatz within the framework of measurement-based VQE [53], and explore the effectivity of having multi-qubit $Z$ rotation gates in the ansatz where the problem Hamiltonian has many-qubit terms.
Another avenue to explore is to ask why it might be the case that this $\mathbb{Z}_{2}$ LGT is suited to VQE. Where fault-tolerant quantum algorithms might succeed and variational approaches might fail is within the circuit unitaries. Fault-tolerant quantum algorithms rely on precisely engineered unitary gates, while VQEs have much redundancy here. Then perhaps the $\mathbb{Z}_{2}$ LGT benefits from being a theory made of simple Pauli operators, that is naturally encoded in terms of qubits. It would be important to extend these studies to other gauge groups, and to further explore the relevance of gauge symmetries. The connection between VQEs and symmetries has been explored, but less so VQEs and gauge symmetries, which could be interesting to further develop.
Acknowledgements.
This work has been supported with funds from the Ministry of Science, Research and Culture of the State of Brandenburg within the Center for Quantum Technology and Applications (CQTA). It also received support from the Federal Ministry for Economic Affairs and Climate Action (project Qompiler, grant No: 01MQ22005A), the DFG (Emmy Noether programme, grant No. 418294583) and the Berlin Quantum Alliance. [Uncaptioned image]

### v3

@@
 In this work, we design a variational quantum eigensolver (VQE) suitable for investigating ground states and static string breaking in a $\mathbb{Z}_{2}$ lattice gauge theory (LGT).
-We consider a two-leg ladder lattice coupled to Kogut-Susskind staggered fermions and verify the results of the VQE simulations using tensor network methods.
+We consider a two-leg ladder lattice with Kogut-Susskind staggered fermions and verify the results of the VQE simulations using tensor network methods.
 We find that for varying Hamiltonian parameter regimes and in the presence of external charges, the VQE is able to arrive at the gauge-invariant ground state without explicitly enforcing gauge invariance through penalty terms.
 Additionally, experiments showing string breaking are performed on IBM’s quantum platform.
-Thus, VQEs are seen to be a promising tool for $\mathbb{Z}_{2}$ LGTs, and could pave the way for studies of other gauge groups.
+Thus, VQEs are seen to be a promising tool for $\mathbb{Z}_{2}$ LGTs, and could serve as a stepping stone toward studies of other gauge groups.
 We find that the scaling of gradients with the number of qubits is favorable for avoiding barren plateaus.
-At the same time, it is not clear how to efficiently simulate the LGT using classical methods.
 Furthermore, strategies that avoid barren plateaus arise naturally as features of LGTs, such as choosing the initialization by setting the Gauss law sector and restricting the Hilbert space to the gauge-invariant subspace.
@@
 At the same time, analytical access in the non-perturbative regime is often not possible.
-One way forward is lattice gauge theory (LGT), in which space-time is discretized on a lattice [1].
+One way forward is lattice gauge theory (LGT), in which space-time is discretized on a lattice [Wilson1974].
 In this framework, matter fields are assigned to the lattice sites, while gauge fields are associated with the connecting links, all while maintaining gauge symmetry.
@@
 Current quantum processors are a long way from being used for realistic simulations of gauge theories.
-So far, successful demonstrations have been carried out for low-dimensional toy models [2, 3, 4, 5, 6, 7] of confinement [8, 9, 10], and scattering dynamics  [11, 12, 13, 14, 15].
-For a review, see  [16, 17, 18].
-In this work, we study a $\mathbb{Z}_{2}$ LGT [19] on a two-leg ladder geometry coupled to Kogut-Susskind fermions [20].
+So far, successful demonstrations have been carried out for low-dimensional toy models [Martinez2016, Atas2021, Atas2022, farrell2024scalable, farrell2024quantum, Klco2018, Klco2019, Kokail2018, meth2025simulating, Cochran2024, Cobos2025] of confinement [Mildenberger2025, Ciavarella2025, alexandrou2025realizingstringbreakingdynamics, Crippa2024a], and scattering dynamics  [Davoudi2024, Zemlevskiy2024, chai_fermionic_2025, chai_resource-efficient_2025].
+For reviews, see, e.g.
+Refs.
+[Banuls2019, Banuls2020, Funcke2023a, Halimeh2025].
+In this work, we study a $\mathbb{Z}_{2}$ LGT [Horn1979] on a two-leg ladder geometry coupled to Kogut-Susskind fermions [Kogut1975].
 The two-leg ladder is an intermediate step between one and two dimensions.
@@
 Using $\mathbb{Z}_{2}$ LGT as a toy model, we investigate static aspects of string breaking as a consequence of confinement.
-When two charges are confined in the $\mathbb{Z}_{2}$ LGT, by Gauss law they are necessarily connected by an electric flux tube.
-The flux tube experiences a linearly rising potential proportional to their distance.
-From a certain distance it becomes energetically favorable to break the flux string, and here this is done by forming an additional pair of charges.
+At large values of the coupling, when two charges are confined in the $\mathbb{Z}_{2}$ LGT, by Gauss law they are necessarily connected by an electric flux tube.
+Increasing their distance leads to a linearly rising potential.
+At large distances, it becomes energetically favorable to break the flux string, and here this is done by forming an additional pair of charges.
+In addition, at intermediate values of the coupling, a roughening transition can occur, which is a phase transition where the confining flux tube becomes rough, i.e., its transverse fluctuations become significant.
+This transition has recently been observed in TN simulations [marcantonio_roughening_2025].
 We apply a variational quantum eigensolver (VQE) to $\mathbb{Z}_{2}$ LGT on the two-leg ladder lattice to recover ground state properties and to explore static string breaking.
@@
 While we do not yet have the quantum hardware to answer this question, recent investigations of variational quantum algorithms have unearthed many issues.
-In particular, there seems to be a connection between the trainability of a VQE (the absence of barren plateaus) and efficient classical simulation [21].
+In particular, there seems to be a connection between the trainability of a VQE (the absence of barren plateaus) and efficient classical simulation [cerezo2024doesprovableabsencebarren].
 Thus, it is important that proposals involving barren-plateau free VQEs consider the scaling with system size and the effects of noise.
 In this work, we find features within the framework of $\mathbb{Z}_{2}$ LGT that are compatible with common strategies for avoiding barren plateaus, all the while being nontrivial to classically simulate.
-Warm starts are one such strategy, where one chooses to initialize close to a solution, where there may be narrow gorges [22].
+Warm starts are one such strategy, where one chooses to initialize close to a solution, where there may be narrow gorges [Puig2025] in the search landscape.
 Analogously, for investigations of LGT, it is necessary to initialize to some Gauss law sector, which thereby restricts the search landscape.
-The search process for a VQE is benefited by constraining to an explorable subspace of Hilbert space [23]; by gauge symmetry for LGTs, this is the subspace of the gauge-invariant, physical, states.
+The search process for a VQE is benefited by constraining to an explorable subspace of Hilbert space [Nakanishi19]; by gauge symmetry for LGTs, this is the subspace of the gauge-invariant, physical, states.
 We draw upon these features using two types of ansatz circuits.
@@
 Results are also obtained using IBM’s quantum hardware, where we perform experiments in the region of the transition into the string-broken phase.
-These findings underscore the usefulness of VQEs in the investigation of LGTs.
+These findings underscore the usefulness and suitability of VQEs in the investigation of LGTs.
 II Background and methods
@@
 On this diagram we have also included an example string configuration, through the solid arrow connecting two charges that have been placed.
-The overlapping circles show the change in sign to indicating the placed charges.
+The overlapping circles show the change in sign, indicating the placed charges.
 The defining feature of gauge theories is invariance under local transformations.
 This leads to the presence of a local conservation law or a continuity equation that determines the interplay between the matter and gauge degrees of freedom.
-In the Hamiltonian formulation using temporal gauge, this results in an additional constraint known as Gauss law.
+In the Hamiltonian formulation using temporal gauge, this results in an additional constraint known as Gauss law [Kogut1975].
 Lattice gauge theories (LGTs) provide a means to probe gauge theories numerically.
@@
 The pure gauge part of the theory is constructed following Ref.
-[19] and the lattice notation for the two-dimensional ladder geometry we consider is illustrated in Fig.
+[Horn1979] and the lattice notation for the two-dimensional ladder geometry we consider is illustrated in Fig.
 1.
@@
 The parameter $\mu>0$ tunes the relative strength between the electric and magnetic energy.
-To couple the theory to fermionic matter, we consider Kogut-Susskind staggered fermions [20].
+To couple the theory to fermionic matter, we consider Kogut-Susskind staggered fermions [Kogut1975].
 This formulation essentially corresponds to separating the spinor components to different lattice sites, resulting in alternating the signs of the masses and charges across the lattice as reflected in the orange and blue coloring in Fig.
@@
 The prefactor represents the different nature of the odd and even lattice sites, and we define the number operator $n_{l}=\phi_{l}^{\dagger}\phi_{l}$.
-Depending on the point of the lattice, the operators acting on the links are replaced by the boundary condition $X_{\bm{l},\bm{l}+\hat{x}}\!\rightarrow 1$.
+Following Ref.
+[alexandrou2025realizingstringbreakingdynamics], we interpret $q_{\bm{l}}=1$ as the absence of a charge, whereas a value of 1-1 corresponds to a static charge at site $\bm{l}$.
+Depending on the point of the lattice, the operators acting on the links are replaced by the boundary condition $X_{\bm{l},\bm{l}+\hat{x}}\!\rightarrow 1$ (see Ref.
+[alexandrou2025realizingstringbreakingdynamics] for details).
 Since $\forall\bm{l}$, $H,G_{\bm{l}}]=0$, $G_{\bm{l}}$ is a local symmetry that decomposes the Hilbert space into separate sectors depending on the charge.
@@
 Gauge-invariant states subject to gauge-invariant operators remain in this physical subspace, i.e., the symmetry remains intact.
-This subspace is still exponentially large, while being non-trivial to split into further gauge-invariant subspaces  [24, 25].
-Incorporating terms proportional to (3) is also a means to introduce infinitely massive static charges to the theory.
+This subspace is still exponentially large, while being non-trivial to split into further gauge-invariant subspaces  [Buividovich_2009, Tagliacozzo2014].
+Incorporating terms proportional to (3) allows for enforcing Gauss law, where infinitely massive static charges can be introduced to the theory by setting the corresponding $q_{\bm{l}}=-1$.
 A consequence of Gauss law is that a pair of charges placed on matter sites in vicinity of each other are necessarily connected by a flux tube or string.
@@
 1, where a configuration of a string is shown between two sites with charges, indicated by the opposite coloring.
-Static string breaking and confinement has also been investigated for $\mathbb{Z}_{2}$ LGT in the following works [26, 27, 28, 10].
+Static string breaking and confinement has also been investigated for $\mathbb{Z}_{2}$ LGT in the following works [Lumia2022Z2, Xu2025, marcantonio_roughening_2025, alexandrou2025realizingstringbreakingdynamics].
 Then, the total Hamiltonian is as follows,
@@
 with $V>0$. Placing charges amounts to adding penalty terms on the appropriate site with the opposite sign.
-To compute expectation values of this Hamiltonian on a quantum computer, we perform the Jordan-Wigner transformation [29] on the matter Hamiltonian (2) that couples to fermionic degrees of freedom.
+To compute expectation values of this Hamiltonian on a quantum computer, we perform the Jordan-Wigner transformation [Hamer1997Series] on the matter Hamiltonian (2) that couples to fermionic degrees of freedom.
+For this we must choose a linear ordering of the fermions.
 The gauge part and penalty terms already consist of Pauli operators.
-For this we must choose a linear ordering of the fermions.
 In Fig.
@@
 One way to reduce the dimensionality of the parameter space is to apply the symmetries of the system.
-Herein lies what enables efficient classical simulation – this reduction tends to identify some polynomially large subspace [21].
+Herein lies what enables efficient classical simulation – this reduction tends to identify some polynomially large subspace [cerezo2024doesprovableabsencebarren].
 As such, it is possible to diagnose barren plateaus from the size of the dynamical Lie algebra (DLA) spanned by the gate generators.
 For the gates of the parametrized ansatz circuit, the DLA generates any of the arbitrary circuits that it can express.
-In [30] it is conjectured that the gradients of the cost function scale inverse-polynomially with the dimension of the DLA.
-However, while a small DLA may avoid barren plateaus, a polynomial-sized DLA implies the problem is classically simulable [31].
-Classical simulation can be performed using techniques such as low-weight Pauli propagation [32, 33].
+In [Larocca2022diagnosingbarren] it is conjectured that the variance of gradients of the cost function scale inverse-polynomially with the dimension of the DLA.
+However, while a small DLA may avoid barren plateaus, a polynomial-sized DLA implies the problem is classically simulable [goh2025liealgebraicclassicalsimulationsquantum].
+Classical simulation can be performed using the DLA [goh2025liealgebraicclassicalsimulationsquantum], or techniques such as low-weight Pauli propagation [schuster2024polynomialtimeclassicalalgorithmnoisy, angrisani2024classicallyestimatingobservablesnoiseless].
 These are important considerations for ansatz design.
@@
 Whereas an ansatz based on unitaries formed of the problem Hamiltonian is known as the Hamiltonian variational ansatz (HVA).
-For the local Hamiltonian $H=H_{K}$, where each $H_{K}$ consists of commuting Pauli operators, the HVA can be written [34, 35] as:
+For the local Hamiltonian $H=H_{K}$, where each $H_{K}$ consists of commuting Pauli operators, the HVA can be written [Wecker2015, Wiersema2020] as:
 […]
-Initialization strategies to avoid barren plateaus have been proposed for both the HEA [36] and HVA [37].
-However, such warm starts and initialization strategies show promise for VQEs, but are not guaranteed to find the global minimum, nor is there a general initialization strategy in subspaces containing the solution.
-This occurs for only a vanishing fraction of problem instances [22].
-In [38] it is proven that local Hamiltonians training finite or logarithmic local-depth circuits (FLDC) will not encounter barren plateaus.
+Initialization strategies to avoid barren plateaus have been proposed for both the HEA [Park:2024rim] and HVA [Park2024hamiltonian].
+However, while warm starts and related initialization strategies show promise for VQEs, they are not guaranteed to find the global minimum.
+Nor does there exist a general initialization strategy in subspaces containing the solution.
+This occurs for only a vanishing fraction of problem instances [Puig2025].
+In [Zhang2024Absence] it is proven that local Hamiltonians training finite or logarithmic local-depth circuits (FLDC) will not encounter barren plateaus.
 Here the two-dimensional Toric code is considered, which is a variant of the pure gauge part of the $\mathbb{Z}_{2}$ LGT.
@@
 Iterating this procedure until convergence, the final MPS provides an approximation for the ground state of the Hamiltonian.
-For our numerical calculations, we use the ITensor library where we run a standard two-site algorithm updating two neighboring tensors together [41, 42].
+For our numerical calculations, we use the ITensor library where we run a standard two-site algorithm updating two neighboring tensors together [Fishman_2022, Fishman2022b].
 This allows for adaptively adjusting the bond dimension during the computation to obtain results with a certain precision.
 A detailed review on MPS methods can be found in Ref.
-[43].
+[Schollwoeck2011].
 Note that in order to apply the MPS approach to our ladder geometry, we have to map the local degrees of freedom sitting on the vertices and links to a linear chain.
@@
 Preserving gauge invariance is also difficult, since errors in the system evolution can push the state outside of the physical Hilbert space.
-One way to enforce and preserve gauge invariance is by constructing a gauge-preserving ansatz [44, 10], as shown in Fig.
+One way to enforce and preserve gauge invariance is by constructing a gauge-preserving ansatz [mazzola2021gauge, alexandrou2025realizingstringbreakingdynamics], as shown in Fig.
 2a.
@@
 2b and we refer to it as ZZ.
-The gates used in this ansatz are inspired by measurement-based quantum computing (MBQC) [45], consisting of single qubit rotation gates, and multi-qubit $Z$ rotations for the entangling layers.
-Such multi-qubit rotations can be implemented in MBQC by the introduction of one ancillary qubit [46].
+The gates used in this ansatz are inspired by measurement-based quantum computing (MBQC) [mbqc_def], consisting of single qubit rotation gates, and multi-qubit $Z$ rotations for the entangling layers.
+Such multi-qubit rotations can be implemented in MBQC by the introduction of one ancillary qubit [qin2024applicability].
 The choice of simpler gates also makes this ansatz more hardware-efficient.
@@
 For the simulations we have averaged over 100 samples and find favorable scaling of the variance up to system sizes of four plaquettes and here the Hamiltonian is the same as in Fig.
+6.
+For both ansatz circuits, we find that when scaling from one to two plaquettes, computing the DLA has too large a computational cost.
+The GI ansatz has similarities with the DLA of the matchgate circuit, which is known to be barren-plateau free and classically simulable [diaz2023showcasingbarrenplateautheory].
+However, the presence of $X$ gates make it so that the size of the DLA for the GI ansatz of one plaquette is $4080$, while the matchgate circuit on the same number of qubits ($8$) has dimension $120$.
+Another example where the scaling of the variance of the cost function changes upon the inclusion of additional unitaries is shown in [Larocca2022diagnosingbarren].
+In order to consider some scaling relation of the DLA, we consider the non-physical scenario of adding fermions to the two-dimensional system one-by-one.
+This amounts to beginning with the smallest LGT, where we have two fermion qubits connected by a link qubit and the system size is increased by adding a fermion qubit alongside a respective link qubit.
+Up to $7$ qubits, the two-leg ladder is the same as the one-dimensional chain, so we cannot use these points to compute the scaling.
+Results for the dimension of the DLA, $\dim\mathfrak{g}$ and scaling is shown in Fig.
 5.
-We also find that for both ansatz circuits, when scaling from one to two plaquettes, computing the DLA has too large a computational cost.
-The GI ansatz has similarities with the DLA of the matchgate circuit, which is known to be barren-plateau free and classically simulable [47].
-However, the presence of $X$ gates make it so that the size of the DLA for the GI ansatz of one plaquette is $4080$, while the matchgate circuit on the same number of qubits ($8$) has dimension $120$.
-Another such example where the scaling of the variance of the cost function changes upon the inclusion of additional unitaries is shown in [30].
+Here we see that the one-dimensional problem of both GI and ZZ ansatz show exponential scaling, and show some data for the two-leg ladder lattice.
+Figure 5: Scaling of the dimension of the DLA.
+For increasing number of qubits of the GI and ZZ ansatz and the one- and two-dimensional LGT.
+Here we include an exponential and polynomial fit for the one-dimensional cases.
+For the GI ansatz the exponential fitting is given by $\dim(\mathfrak{g})\approx$3.49\text{\times}{10}^{-1}$2^{1.50n}$, where $n$ is the number of qubits, and the mean squared error (MSE) $10.56$, while the polynomial fit is $\dim(\mathfrak{g})\approx$5.71\text{\times}{10}^{-7}$n^{10.33}$ with MSE $8550.98$.
+Whereas for the ZZ ansatz the exponential fit is $\dim(\mathfrak{g})\approx$5.47\text{\times}{10}^{-1}$2^{1.98n}$ with MSE $1.63$, and then the polynomial fit is $\dim(\mathfrak{g})\approx$1.04\text{\times}{10}^{-3}$n^{8.17}$ with MSE $258.65$.
+Up to $7$ qubits, the one- and two-dimensional cases are the same, as the 8th qubit introduces a plaquette term.
+With this additional term, the DLA is larger than for the one-dimensional case.
 Figure 6: Performance of the VQE.
@@
 For each instance we have taken the average of 2020 runs and used the SPSA optimizer.
-In the noiseless case, low-weight Pauli propagation (PP) is a useful technique for computing expectation values where the problem is efficiently classically simulable [33].
-However, for these simulations we find large errors and ineffective weight truncation.
-Importantly, these ansatz circuits use non-Clifford rotation gates, so that their entangling layers do not meet the criteria to use this method, but all the while evading barren-plateaus.
 The LGT framework provides natural parallels with the features of VQEs that escape barren plateaus.
@@
 Gauss law also acts as a signal to when the simulation may be trapped in local minima, since Gauss law is violated for these states.
-For VQEs to work, the optimization uses symmetries of the problem, while here we have the gauge-invariant subspace.
-Furthermore, a problem is classically simulable if this subspace is polynomially sized, but we know the overall gauge-invariant subspace to be exponential in size [24, 25].
-For the results, classical simulation of the VQE was performed using Pennylane [48], and to interact with IBM’s quantum platform, we have used Qiskit [49].
+For VQEs to work, the optimization leverages symmetries of the problem, and here this is the gauge-invariant subspace.
+Furthermore, a problem is classically simulable if this subspace is polynomially sized, but we know the overall gauge-invariant subspace to be exponential in size [Buividovich_2009, Tagliacozzo2014].
+For the results, classical simulation of the VQE was performed using Pennylane [bergholm2022pennylaneautomaticdifferentiationhybrid], and to interact with IBM’s quantum platform, we have used Qiskit [javadiabhari2024quantumcomputingqiskit].
 III.2 Ground state estimation
@@
 We show results for varying the numbers of shots, system size and numbers of layers for an instance of a Hamiltonian configuration in Fig.
-5.
+6.
 Here we use $J=3$, $m=1$, $\mu=2$, run simulations up to $3$ layers (see Fig.
@@
 In Fig.
-6, we plot the optimization to find the ground state energy of the Hamiltonian in Eq.
+7, we plot the optimization to find the ground state energy of the Hamiltonian in Eq.
 (4) with $J=5$, $m=1$, $\mu=2$ and charges placed at $(0,0)$ and $(0,2)$.
 This case is interesting as the corresponding TN calculation is stuck in a local minima, with this value plotted as a horizontal line.
-Thus the TN calculation for comparison requires the addition of penalty terms to the Hamiltonian, while this is not true for the VQE using either GI or ZZ ansatz.
-This can occur since TN approaches tend to be biased towards weakly correlated phases.
 Figure 7: VQE optimization of the ground state.
@@
 Without the addition of penalty terms, the VQE is successful in converging to the ground state.
-In the inset we show results also for fixed numbers of shots $[500,1000,5000,10000]$ averaged over 1515 runs.
-In order to consider fixed numbers of shots, we optimize using SPSA.
+In the inset we show results also for fixed numbers of shots $[500,1000,5000,10000]$ averaged over $15$ runs.
+In the software, rather optimizing with SPSA enables us to consider fixed numbers of shots.
 It seems the final energy does not depend on the number of shots.
@@
 The results for string breaking are captured by the static potential $V(d)$, where $V(d)=E_{0}-E_{c}$, and $E_{c}$ is the ground state with charges in Fig.
-7.
-The lattice spacing $d$ is taken with reference to the fermion at $(0,0)$, so $d$ counts the number of links between this and the other available fermion.
-This means that for $d=1,2,3$ there are two fermion sites and $d=4$ contains the furthest fermion from $(0,0)$ at $(3,1)$.
-The oscillations in $V(d)$ for the two fermions per fixed $d$ is due to the boundary conditions and superpositions of string configurations, where this effect is heightened at the critical $d=2$.
+8.
+The lattice distance $d$ is taken with reference to the fermion at $(0,0)$, so $d$ counts the number of links between this and the other available fermion.
+This means that for distances $d=1,2,3$ there are two possible fermion sites to place the static charge, whereas for $d=4$ there the only choice is site $(3,1)$.
+The different values in $V(d)$ for fixed distances $0<d<4$ are due to the different possible positions of the static charges and due to the corresponding boundary conditions, where this effect is heightened at the critical $d=2$.
 Here the charge at $(2,0)$ still has the presence of a flux string, while for $(1,1)$ we find the string vanishes.
@@
 This behavior is shown in the lattice configuration plots of Fig.
-7.
+8.
 Here we have measured the observable $\langle Z\rangle$ on the matter qubits and $\langle X\rangle$ on the link qubits.
@@
 The resultant 141 angles are then used to prepare the state on ibm_marrakesh, where the transpiled circuit has a depth of 525 and a $CZ$ count of 436.
-We use the default IBM error mitigation of zero-noise extrapolation (ZNE) and Twirled Readout Error eXtinction (TREX) [50], and take a plug-and-play approach to the experiments, where we choose the best of 5 runs.
+We use the default IBM error mitigation of zero-noise extrapolation (ZNE) and Twirled Readout Error eXtinction (TREX) [vandenBerg2022Modelfree], and take a plug-and-play approach to the experiments, where we choose the best of 5 runs.
 For the complicated circuits it is remarkable that the qualitative behavior of the VQE can be seen, but the results indicate that we should consider smaller system sizes.
@@
 The experiments are of a single run only (as though we are performing the last step of the VQE optimisation on the quantum device) and the results for the resulting ground state energy $E$ are plotted in Fig.
-8, where we have used the default ZNE and TREX error mitigation.
+9, where we have used the default ZNE and TREX error mitigation.
 For this case there is strong agreement of the VQE with three layers of the ZZ ansatz and the TN simulations.
@@
 The GI ansatz has comparable performance with fewer numbers of layers, (see Fig.
-5) so it would be interesting to perform similar analysis with one layer of the GI ansatz.
+6) so it would be interesting to perform similar analysis with one layer of the GI ansatz.
 The degree to which the ZZ ansatz is hardware efficient can also be determined in this way.
@@
 In doing so, we find that $\mathbb{Z}_{2}$ LGTs have features that suit investigation using VQEs, since they avoid barren plateaus and can be nontrivial to classically simulate.
-One way to avoid barren plateaus is to use smart initialization strategies [22]; however a general strategy to identify these or to avoid local minima is not known.
+One way to avoid barren plateaus is to use smart initialization strategies [Puig2025]; however a general strategy to identify these or to avoid local minima is not known.
 For LGTs on the other hand, we always initialize in some Gauss law sector and would know beforehand which to initialize in.
@@
 When the subspace spanned by the ansatz is polynomial in size, the problem can be classically simulated.
-All the while, the full space of gauge-invariant states is exponentially large [24, 25].
+All the while, the full space of gauge-invariant states is exponentially large [Buividovich_2009, Tagliacozzo2014].
 In this work we present two ansatz circuits, one that is constructed of gauge-invariant Hamiltonian terms, this we refer to as GI, and another that is inspired by MBQC that we name ZZ, as it is composed of multi-qubit $Z$ rotation gates.
@@
 This suggests that problems of LGTs could be useful for testing VQEs at scale.
-Fundamentally, the VQE could never be performed in ideal conditions, so considering the effects of scaling with system size, number of shots and incorporating the effects of noise is vital for making such a determination.
+Fundamentally, VQE would never be implemented in ideal conditions, so considering the effects of scaling with system size, number of shots and incorporating the effects of noise is vital for making such a determination.
 We find that even with low numbers of shots $\sim 500$, convergence to the ground state is close and the resulting state is gauge-invariant.
@@
 In future work we would look to perform more experiments using quantum hardware.
-For smaller system sizes, it would be useful to perform some rounds of optimization in the VQE to see whether the resulting state remains gauge-invariant.
-This can be extended to observations of the string-breaking transition, where it would be interesting to see if this persists.
-We would also be interested to see the effect of small penalty terms on the cost landscape, since small terms in a region $\sim J$ can benefit the convergence of the $ZZ$ ansatz.
+For smaller system sizes, it would be useful to perform some rounds of optimization in the VQE to see whether the resulting state remains gauge-invariant, and extending to observations of the string-breaking transition, where it would be interesting to see if this persists.
+We are also interested in the effects of small penalty terms on the cost landscape, since small terms in a region $\sim J$ can benefit the convergence of the $ZZ$ ansatz.
 Here one can also investigate the effect of small penalty terms in mitigating noise that may push the state out of the gauge-invariant subspace.
-We would also consider larger ladder lattices to mitigate the boundary effects.
-Some preliminary studies involving square lattices have also been performed, and show that we can extend this approach.
+Larger ladder lattice sizes would mitigate the boundary effects and we have performed preliminary studies involving square lattices, showing that we can extend this approach.
 The behavior of the GI ansatz using current hardware versus the more hardware-efficient ZZ ansatz can also be analyzed, which would indicate the practicality of scaling such a VQE in the NISQ era.
@@
 It would also be beneficial to extend these studies to other platforms too.
-For instance the long-range multi-qubit interaction gates of the two ansatz may better suit ion-trap [51] or neutral atom based [52] quantum hardware.
-It is also interesting to explore applying the ZZ ansatz within the framework of measurement-based VQE [53], and explore the effectivity of having multi-qubit $Z$ rotation gates in the ansatz where the problem Hamiltonian has many-qubit terms.
+For instance the long-range multi-qubit interaction gates of the two ansatz may better suit ion-trap [Molmer2000] or neutral atom based [baker2021exploitinglongdistanceinteractionstolerating] quantum hardware.
+It is also interesting to explore applying the ZZ ansatz within the framework of measurement-based VQE [Ferguson2021VQE], and explore the effectivity of having multi-qubit $Z$ rotation gates in the ansatz where the problem Hamiltonian has many-qubit terms.
 Another avenue to explore is to ask why it might be the case that this $\mathbb{Z}_{2}$ LGT is suited to VQE.